“All changes, even the most longed for, have their melancholy; for what we leave behind us is a part of ourselves; we must die to one life before we can enter another.”

Anatole France

‘The moment a man starts to behave ridiculous to a woman you know he is serious at it.

Colette

1

Quote to be quoted

This is an example of a normal chapter page.

2

Enjoy the whitespace

This is an example of a chapter without quotes.

More space for tocs and bibs

This is an example of a non-numbered page.

Who’s gonna U:~ tonight?

The Cars

Do not trust everyone

This is to show there is not much space in the star version chapters.